OpenStudy (anonymous):

Does anyone want to help me with six questions on Graphing Linear and Absolute Value Equations and Inequalities?

OpenStudy (anonymous):

@jim_thompson5910 could you help me?

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

name one

OpenStudy (anonymous):

the most basic one is \[y=|x|\] which looks like this 

OpenStudy (anonymous):

here is the attachment of the assignment

OpenStudy (anonymous):

you can plot some points if you like, or maybe we can do the following take the graph i wrote above \[y=|x|\] then for \[y=|x-2|\] you shift it to the right two units, then for \[y=-\frac{1}{5}|x-2|\] flip it up side down and make the slops of the lines \(\frac{1}{5}\) and \(-\frac{1}{5}\) instead of \(-1\) and \(1\) then shift it up 3 units to get \[y=-\frac{1}{5}|x-2|+3\]

OpenStudy (anonymous):

here is a picture from wolfram http://www.wolframalpha.com/input/?i=y%3D-1%2F5 |x-2|%2B3

OpenStudy (anonymous):

we could also do this make a point at \((2,3)\) which is the highest point on the graph

OpenStudy (anonymous):

then plot a couple other well chosen points, say \(x=7\) gives you \[y=-\frac{1}{5}|7-2|+3=-1+3=2\] so plot \((7,2)\)

OpenStudy (anonymous):

and \(x=-3\) give you \[y=-\frac{1}{5}|-3-2|+3=-1+3=2\] so plot \((-3,2)\) connect the points \((2,3)\) and \((-3,2)\) with a straight line connect the points \((2,3)\) and \((7,2)\) with another straight line you want \[y\geq -\frac{1}{5}|x-2|+3\] so shade everything above those two lines

OpenStudy (anonymous):

could you do this on the assignment I attached?